### v2

# Data-Driven Learnability Transition of Measurement-Induced Entanglement

Abstract
Measurement-induced entanglement (MIE) captures how local measurements generate long-range quantum correlations and drive dynamical phase transitions in many-body systems. Yet estimating MIE experimentally remains challenging: direct evaluation requires extensive post-selection over measurement outcomes, raising the question of whether MIE is accessible with only polynomial resources. We address this challenge by reframing MIE detection as a data-driven learning problem that assumes no prior knowledge of state preparation. Using measurement records alone, we train a neural network in a self-supervised manner to predict the uncertainty metric for MIE—the gap between upper and lower bounds of the average post-measurement bipartite entanglement. Applied to random circuits with one-dimensional all-to-all connectivity and two-dimensional nearest-neighbor coupling, our method reveals a learnability transition with increasing circuit depth: below a threshold, the uncertainty is small and decreases with polynomial measurement data and model parameters, while above it the uncertainty remains large despite increasing resources. We further verify this transition experimentally on current noisy quantum devices, demonstrating its robustness to realistic noise. These results highlight the power of data-driven approaches for learning MIE and delineate the practical limits of its classical learnability.
Quantum entanglement is a central resource of quantum information science, enabling advantages in computation, communication, and sensing [1, 2, 3, 4, 5]. Importantly, entanglement need not arise solely from coherent unitary evolution: suitably chosen and adaptively processed measurements can also generate nonlocal correlations in many-body systems [6, 7, 8, 9]. This measurement-induced entanglement (MIE) underlies measurement-based quantum computation [10], enables rapid preparation of long-range entangled states [11, 12, 13, 14, 15], and gives rise to novel non-equilibrium phases of matter [16, 17, 18, 19, 20]. In particular, in hybrid unitary–measurement dynamics, competition between scrambling and projective measurements produces a measurement-induced phase transition (MIPT), across which the scaling of MIE changes from volume law to area law [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
Despite its conceptual and practical significance, directly characterizing MIE in experiments remains notoriously challenging [38]. The central obstacle is post-selection: probing properties of a state conditioned on a specific measurement outcome requires repeating the experiment until that outcome reoccurs, an effort that grows exponentially with the number of measurements by Born’s rule. Several scalable diagnostics have been proposed to circumvent this bottleneck, including purification of an entangled reference qubit [39, 40, 41, 42, 43], learnability of conserved quantities [44] or of the pre-measurement state [45], cross-entropy benchmarks [46, 47, 48] and other machine-learning proxies [49, 50, 51]. Although these proxies successfully reflect the distinct behavior of MIE in different regimes and help reveal its critical behavior, they remain indirect witnesses rather than quantitative estimators of MIE itself. Recent works have made progress showing promise and limits [52, 53]. On the one hand, MIE can in principle be estimated without post-selection by leveraging quantum-classical correlations, but this requires priori knowledge of the underlying quantum dynamics and the accuracy of estimation hinges on the fidelity of classical simulations. On the other hand, Ref. [53] proves that, without any knowledge of the pre-measurement state, no learning protocol can extract properties beyond ensemble averages using only a polynomial number of measurement shots. The practical limits of experimentally detectable MIE thus remain an open question.
Motivated by the recent success of data-driven quantum learning methods based solely on measurement records [54, 55, 56], we ask in this Letter: under what conditions can MIE be learned with polynomial resources from data alone—i.e., without post-selection and without any prior knowledge of state preparation? To address this question, we study the entanglement generated between two distant qubits $A$ and $B$ after all other qubits in a many-body state are measured. Using measurement outcomes only, we train a transformer-based neural network in a self-supervised manner to estimate the post-measurement state on $AB\equiv A\cup B$ conditioned on a given measurement outcome. Instead of reconstructing MIE directly, we estimate the total entanglement entropy of $AB$, which we use as an uncertainty metric quantifying the learnability of MIE. We focus on two settings—random one-dimensional (1D) all-to-all circuits and two-dimensional (2D) nearest-neighbor circuits—and in both we observe a clear learnability transition with increasing circuit depth. In the learnable phase, the uncertainty decreases with additional measurement data and model parameters, reflecting increasing time and space complexity. In the unlearnable phase, the uncertainty saturates or even grows. We further probe this learnability transition in the presence of noise through simulations and experiments on IBM QPU ibm_marrakesh, finding that the transition remains observable under realistic noise. Overall, our results demonstrate the power of data-driven methods to extract informative features of the post-measurement state that reveal MIE, while simultaneously clarifying the fundamental limitations of learning MIE from measurement data alone. We note that a recent study [57] investigated MIE learnability for GHZ and cluster states using a similar data-driven approach, where signatures of a learnability transition were observed for the latter upon tuning the measurement direction. In contrast, our work addresses generic states generated by random quantum circuits and provides an explicit characterization of the emergence of distinct learnable and unlearnable regimes.
[…]
$S_{A,m}^{\text{QC}}\geq S_{A,m}\geq S_{A,m}^{\text{QC}}-S_{AB,m}^{\text{QC}},$ (1)
which utilizes $S_{AB,m}\geq 0$. Thus, the true entanglement $S_{A,m}$ lies within an interval of width $S_{AB,m}^{\text{QC}}$ for each outcome $m$. Nevertheless, directly computing $S_{AB,m}^{\text{QC}}$ requires complete knowledge of $\sigma_{AB,m}$. To access this quantity, we employ classical shadow tomography [58]. Qubits $A$ and $B$ are measured in random single-qubit Pauli basis, producing classical snapshots $\sigma_{AB,m}^{s}$ that are an unbiased estimators satisfying $\mathbb{E}_{s}[\sigma_{AB,m}^{s}]=\sigma_{AB,m}$, where $\mathbb{E}_{s}$ denotes averaging over different snapshots. We then define
$\Delta\equiv\mathbb{E}_{m}\mathbb{E}_{s}\left[S_{AB,m}^{\text{SC}}\right]=\mathbb{E}_{m}\left[S_{AB,m}^{\text{QC}}\right],$ (2)
[…]
Machine-learning model—Rather than estimating $\rho_{AB,m}$ by classically simulating the state-preparation dynamics, we learn it directly from the measurement data. We parameterize $\rho_{AB,m}$ with a neural network and train it to approximate the true post-measurement state $\sigma_{AB,m}$. To construct the loss function, we use the inequality $\mathbb{E}_{m}[\text{Tr}(\sigma_{AB,m}-\rho_{AB,m})^{2}]\geq 0$, which implies
$\mathbb{E}_{m}\mathbb{E}_{s}\left[2\text{Tr}(\rho_{AB,m}\sigma_{AB,m}^{s})-\text{Tr}(\rho_{AB,m}^{2})\right]$ (3)
$\leq\mathbb{E}_{m}\left[\text{Tr}\sigma_{AB,m}^{2}\right]\leq 1.$
The first equality saturates if and only if $\rho_{AB,m}=\sigma_{AB,m}$ for all $m$, and the second becomes an equality when $\sigma_{AB,m}$ is pure. This motivates the loss function:
$\mathcal{L}(\theta)=-\mathbb{E}_{m}\mathbb{E}_{s}\left[2\text{Tr}(\rho_{AB,m}\sigma_{AB,m}^{s})-\text{Tr}(\rho_{AB,m}^{2})\right],$ (4)
where $\theta$ denotes the model parameters. We adopt this loss rather than directly targeting $\Delta$ because $\text{Tr}(\rho_{AB,m}\sigma_{AB,m}^{s})$ is bounded, while $S_{AB,m}^{\text{SC}}$ exhibits unbounded fluctuations [52]. Alternative loss functions were also examined; detailed comparisons are provided in the Supplemental Material [59]. Importantly, this approach uses only experimentally accessible data: the labels $\sigma_{AB,m}^{s}$ are directly reconstructed from measurements on $A$ and $B$, enabling fully self-supervised training.
Figure 1: Setup and learnability transition. (a) Schematic workflow. Given a quantum state, random single-qubit rotations $U_{A}^{s}$ and $U_{B}^{s}$ are applied to qubits $A$ and $B$, after which all qubits are measured projectively in the computational basis. Measurement outcomes on all qubits except AA and $B$ are fed into a transformer encoder that outputs $\rho_{AB,m}$, while the outcomes on $A$ and $B$ are used to construct the classical snapshot $\sigma_{AB,m}^{s}$. The loss $\mathcal{L}(\theta)$ and uncertainty $\Delta$ are computed from these quantities. Notably, no prior knowledge of the state-preparation procedure is required. (b) Structure of the 1D random all-to-all circuit. (c) $\mathcal{L}(\theta)$ as a function of circuit depth $t$ for different system size $L$. Error bars denote the standard error over $M=50$ different circuit realizations. (d) $\Delta$ as a function of $t$ for different $L$. The inset shows the MIE between $A$ and $B$. Each point is averaged over $5\times 10^{4}$ circuit realizations; the associated standard error is smaller than the symbol size.
For the model architecture, we employ a transformer-based encoder consisting of multi-head self-attention layers and feed-forward blocks [60, 61]. Transformers efficiently capture long-range correlations, making them well-suited for modeling the nonlocal dependence of $\rho_{AB,m}$ on global measurement patterns. Each measurement outcome $m$, represented as a binary sequence of $\{+1,-1\}$, is augmented with a $\text{CLS}]$ token at the beginning and encoded into high-dimensional representations. A specialized density matrix head then maps the $\text{CLS}]$ token’s hidden state to a $4\times 4$ complex-valued density matrix, with hermiticity, positive semi-definiteness, and unit trace constraints automatically satisfied by construction. A schematic workflow is shown in Fig. 1(a), with architectural details provided in [59]. For a fixed quantum state, the training set contains $N_{m}$ measurement shots, which defines the quantum time complexity, while the model contains $N_{p}$ trainable parameters, reflecting the classical space complexity. Practical classical learnability without post-selection is feasible only if $N_{m}$ and $N_{p}$ do not grow exponentially with $L$. After training, we evaluate the learned estimator using an independent set of $N_{e}$ shots to compute $\Delta$ via Eq. (2).
Learnability transitions—We first consider states generated by a 1D random all-to-all circuit, as illustrated in Fig. 1(b). At each discrete time step $\delta t$, we apply $L\delta t$ Haar-random two-qubit gates between uniformly chosen pairs of qubits. For each depth $t$, we average over $M=50$ independently sampled circuit instances to capture average-case behavior. All training data are obtained from exact classical simulations of the circuit dynamics. We first train on a system of size $L=20$ using $N_{p}=7\times 10^{4}$ model parameters and $N_{m}=8\times 10^{4}$ measurement shots, both well within the polynomial-resource regime $N_{p},N_{m}\ll 2^{L}$ for learning process. Additional hyperparameters are provided in [59]. Fig. 1(c) shows that the training loss $\mathcal{L}(\theta)$ converges to its lower bound 1-1 for shallow depths. As $t$ increases, the loss rises, indicating that the learned estimator $\rho_{AB,m}$ fails to reproduce the true post-measurement state. To quantify MIE uncertainty, we evaluate $\Delta$ using $N_{e}=5\times 10^{4}$ independent samples. As shown in Fig. 1(d), $\Delta$ increases with $t$ and saturates near $2\log(2)$. This saturation suggests that at large $t$, the model outputs a maximally mixed state regardless of the measurement outcomes. Alternative loss functions may encourage pure-state predictions, but at large $t$ the predicted state becomes nearly orthogonal to the true one, still producing large $\Delta$ [59].
This learnability transition directly reflects an intrinsic computational phase transition in random quantum circuits, closely related to a finite-depth many-body teleportation transition [62]. As shown in the inset of Fig. 1(d), the MIE between $A$ and $B$ exhibits a finite-depth transition. Below a critical depth, the dynamics remain in a non-teleporting regime, correlations are essentially local, MIE vanishes as the system size becomes larger, and efficient classical algorithms exist for approximate random-circuit sampling and the dynamics exhibit low effective complexity [63]. In this regime, states are expected to admit efficient learning [64], implying that suitably designed neural networks can achieve accurate estimation with polynomial data and parameters resources [65]—the learnable phase. Beyond the critical depth, the circuit enters a teleporting regime where long-range entanglement emerges. Efficient classical simulation is believed to fail here, rendering average-case random-circuit sampling computationally intractable and establishing the basis of quantum computational advantage [66]. Therefore, polynomial-resource machine-learning models are not expected to succeed—defining the unlearnable phase. From this perspective, the learnability transition provides a direct machine-learning probe of computational complexity, faithfully tracking the underlying computational phase boundary. To demonstrate the finite-depth nature of this transition, we compare the results across different system sizes $L=24$ using the same $N_{p}$ and $N_{m}$. At small $t$, $\Delta$ for $L=24$ is comparable to or smaller than that for $L=20$. As $t$ increases, the curves cross, indicating that the transition occurs at a finite depth. Accurately determining the critical depth would require simulations at larger $L$ to minimize finite-size effects.
Figure 2: Learnability transition in 1D random all-to-all circuits for $L=20$. (a) $\Delta$ versus depth $t$ for varying $N_{m}$ with $N_{p}=7\times 10^{4}$ fixed. (b) $\Delta$ versus depth $t$ for varying $N_{p}$ with $N_{m}=8\times 10^{4}$ fixed. (c,d) $\Delta$ as a function of $N_{m}$ and $N_{p}$ for representative depths $t=1.25$ and $t=3.5$, respectively.
To further characterize the two phases, we fix $L=20$ and vary $N_{p}$ and $N_{m}$. Detailed model architectures corresponding to different $N_{p}$ are provided in [59]. Throughout we ensure $N_{p},N_{m}\ll 2^{L}$. Fixing $N_{p}=7\times 10^{4}$ and varying $N_{m}$, Fig. 2(a) shows that at small $t$, $\Delta$ decreases as data $N_{m}$ increases, indicating that additional samples improve the model’s ability to extract MIE. In contrast, $\Delta$ saturates at large $t$, suggesting the presence of an unlearnable phase where the polynomial of many samples does not lead to significant improvement. Fixing $N_{m}=5\times 10^{4}$ and varying $N_{p}$, Fig. 2(b) shows that increasing $N_{p}$ improves performance at small $t$, while $\Delta$ saturates or even grows at large $t$. For excessively large $N_{p}$, $\Delta$ increases for all depths due to overfitting from insufficient training data. The combined dependence on $N_{p}$ and $N_{m}$ is shown in Fig. 2(c) and Fig. 2(d) for $\Delta(t=1.25)$ and $\Delta(t=3.5)$, respectively. At $t=1.25$, $\Delta$ clearly decreases with both resources (except in the strongly overparameterized regime where overfitting causes $\Delta$ to grow). By contrast, at $t=3.5$, $\Delta$ rapidly saturates and even grows despite the same increase in $N_{p}$ and $N_{m}$, thus sharply distinguishing the learnable and unlearnable phases.
Figure 3: Learnability transition in 2D nearest-neighbor circuits for $L=5\times 5$. (a) $\Delta$ versus $t$ for varying $N_{m}$ with $N_{p}=7\times 10^{4}$ fixed. (b) $\Delta$ versus $t$ for varying $N_{p}$ with $N_{m}=8\times 10^{4}$ fixed. (c,d) $\Delta$ as a function of $N_{m}$ and $N_{p}$ at representative depths $t=1.6$ and $t=6.4$, respectively.
We next examine 2D circuit with nearest-neighbor couplings. We consider a square lattice with periodic boundary conditions, where at each step $L\delta t$ two-qubit gates act only on adjacent qubit pairs. This setup is also known to exhibit a finite-depth computational transition [63, 67, 68, 69, 70]. For a system of size $L=5\times 5$, varying $N_{p}$ and $N_{m}$ as in the 1D case, we observe the same qualitative behavior: for small $t$, $\Delta$ decreases markedly as either resource increases, indicating a learnable phase in which additional data or model capacity improves the MIE estimate; while at large $t$, $\Delta$ saturates or grows with increasing resources, signaling the unlearnable phase, as shown in Fig. 3(a,b). The dependence of $\Delta(t=1.6)$ and $\Delta(t=6.4)$ on different combinations of $N_{p}$ and $N_{m}$ [Fig. 3(c,d)] mirrors the 1D case, reinforcing the presence of distinct learnable and unlearnable phases.
Figure 4: Learnability transition in 1D all-to-all circuits with noise. (a,b) $\mathcal{L}(\theta)$ and $\Delta$ for $L=16$, obtained from noisy classical simulations using the Qiskit noise snapshot of the IBM QPU ibm_brisbane. Here $N_{m}=2\times 10^{4}$, $N_{p}=2\times 10^{4}$, $N_{e}=5\times 10^{4}$, and $M=50$. (c,d) $\mathcal{L}(\theta)$ and $\Delta$ for $L=20$, obtained from experiments on the IBM QPU ibm_marrakesh. Here $N_{m}=4\times 10^{4}$, $N_{p}=7\times 10^{4}$, $N_{e}=5\times 10^{3}$, and $M=5$.
Noisy device—The preceding analysis assumes ideal, noiseless circuit dynamics. However, in realistic experiments, noise introduces two key effects. First, each nominally identical state-preparation run experiences a different and unknown noise realization, making it harder for the model to extract consistent information from measurement records. Second, noise renders the post-measurement state $\sigma_{AB,m}$ mixed rather than pure, leading to $\mathbb{E}_{m}[S_{AB,m})]>0$. As discussed above, this finite entropy constitutes an intrinsic contribution to the uncertainty of MIE and cannot be eliminated by improving model accuracy. To examine these effects, we combine noisy numerical simulations with direct experiments on current superconducting quantum processors. As a first step, we adopt a noise model based on the Qiskit snapshot of the IBM QPU ibm_brisbane backend. For a 1D all-to-all circuit with $L=16$, Fig. 4(a,b) show that the learnability transition remains visible, although $\Delta$ is systematically larger than that in the ideal case even at small depths $t$. This offset reflects the intrinsic uncertainty induced by noise. Due to the computational cost of noisy simulation at larger scales, we turn to direct experiments on the IBM QPU ibm_marrakesh for $L=20$ [71]. Constrained by available experimental resources, we use $N_{m}=4\times 10^{4}$, $N_{e}=5\times 10^{3}$ and $M=5$. The resulting data, Fig. 4(c,d), show that $\Delta$ increases with $t$ and eventually saturates near $2\log(2)$, demonstrating that the learnability transition persists on current noisy hardware. It is also noteworthy that $\Delta$ at small $t$ is lower than in the ibm_brisbane simulations, consistent with the lower noise level of ibm_marrakesh and the correspondingly reduced intrinsic uncertainty.
Discussions—Our results have several implications for extracting MIE in practice and for understanding fundamental limits of quantum learning. In the shallow-depth regime, the observed learnability indicates that neural networks trained solely on measurement data can serve as surrogates for classical simulation whenever the latter remains tractable. This demonstrates the possibility of deploying data-driven neural network decoders instead of simulation-based decoders for quantum error correction [72, 73, 74, 75, 76]. By contrast, the loss of learnability at large depth provides evidence for quantum advantage: when classical simulation becomes intractable, machine-learning approaches likewise fail to recover MIE with polynomial resources. This further suggests that in deep circuits within our setup, not only is substantial MIE generated, but crucially, this entanglement remains undetectable. Harnessing MIE in such regimes—rather than merely witnessing its presence—thus remains an important outstanding challenge.
Our work also suggests several interesting directions for future investigation. The present resource analysis accounts for quantum time complexity via $N_{m}$ and classical space complexity through $N_{p}$, but does not incorporate the classical time complexity of training itself. Including training overhead would sharpen the boundary of efficient MIE extraction and is essential for a complete complexity-theoretic characterization. In addition, our study has focused on states generated by random circuits, which represent an average-case scenario. Since physically motivated states—e.g., ground states of interacting Hamiltonians or states with specific symmetries—typically exhibit structure absent in random ensembles, exploring learnability transitions in such settings is a natural and important next step. We leave this to future work.
Acknowledgements.
Acknowledgments—We acknowledge Yi-Zhuang You and Pengfei Zhang for valuable discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission.